# Spatially-Augmented Sequence-to-Sequence Neural Diarization for Meetings

### v1

@@ 4.1 DOA Analysis
It is worth noting that the original SRP-DNN was developed for the 12-channel LOCATA dataset [36] and thus not directly applicable to the 8-channel circular array of AliMeeting. To address this, we retrain the CRNN component of SRP-DNN under the AliMeeting configuration. Fig. 2 further illustrates DOA estimation results on both simulated LibriSpeech mixtures and real AliMeeting data. We can see that, simulated data exhibit negligible DOA errors, while in real meetings, azimuth estimates provide clear discrimination across different speakers. Hence, we only use azimuth as the DOA feature for integration into the diarization system. Moreover, since SRP-DNN’s IDL-based detection tracks at most two speakers per frame, we analyze the AliMeeting training set and find that only 5.98% of the duration involves more than two simultaneous speakers, confirming the assumption’s practicality in real meetings.
@@ 5 Conclusions
This work proposes SA-S2SND, a novel diarization system that integrates direction-of-arrival (DOA) cues into a sequence-to-sequence neural diarization. By incorporating DOA features and a staged training strategy, the model unifies both single- and multi-channel inputs, supporting both online and offline inference. Evaluated on the AliMeeting test set, SA-S2SND achieves significant reductions, particularly when enhanced with channel attention. Future work will focus on improving multi-speaker DOA robustness, exploring joint data training strategies, and further enhancing overall system performance.

### v2

@@ 4.1 DOA Analysis
It is worth noting that the original SRP-DNN was developed for the 12-channel LOCATA dataset [locata] and thus not directly applicable to the 8-channel circular array of AliMeeting. To address this, we retrain the CRNN component of SRP-DNN under the AliMeeting configuration. Fig. 2 illustrates DOA estimation results on both simulated LibriSpeech mixtures and real AliMeeting recordings. The simulated data show negligible DOA errors, while real meeting scenarios exhibit clearly separated azimuth estimates across different speakers. In fact, meeting participants are typically seated, resulting in limited elevation variation, while azimuth provides clear and discriminative spatial cues (Fig.  2(b)). Therefore, azimuth is used as the primary spatial feature. Notably, SA-S2SND is not limited to azimuth-only inputs, and elevation can be incorporated without architectural changes when needed.
Furthermore, SRP-DNN’s IDL-based detection tracks at most two speakers per frame, we believe it is reasonable in practice, since frames with more than two active speakers are rare—only 6% in AliMeeting, 0.5% in AISHELL-4 and 3% in AMI (see Table 1), and thus have a negligible impact on overall diarization performance.
@@ 4.2 Results on Alimeeting Dataset
Table 3: Comparisons of our methods with others on the Alimeeting test sets (offline).
[b] Methods DER (%) Offline    Diaper [landini2024diaper] 20.70    EEND-EDA  [Broughton25] 12.30    Pyannote.audio  [plaquet2023powerset] 15.20    EEND-M2F  [harkonen2024eend] 13.20    EEND-TA + FT  [Broughton25] 11.41    WavLM-Large  [han2025efficient] $\dagger$} 10.80    S2SND (E6 in Table 2) 11.33    SA-S2SND (E7 in Table 2) 10.40
  • $\dagger$}
    State-of-the-art by submission.
4.3 Comparison with Other Existing Methods
Table  3 compares our proposed methods with the previous results on the Alimeeting dataset. In the offline scenario, our proposed SA-S2SND obtain the lowest DERs of 10.40%. The WavLM-large method employs a pretrained WavLM with 63.3M parameters and uses a 16s window for both training and inference (DER=15.1% for 8s window). Despite using a shorter 8s window and without pretraining models, our method achieves state-of-the-art performance.
@@ 5 Conclusions
This work proposes SA-S2SND, a novel spatially-augmented diarization system that integrates direction-of-arrival (DOA) cues into a sequence-to-sequence neural diarization. By incorporating DOA features and a staged training strategy, the model unifies both single- and multi-channel inputs, supporting both online and offline inference. Evaluated on the AliMeeting test set, SA-S2SND achieves significant reductions, particularly when enhanced with channel attention. Future work will focus on improving multi-speaker DOA robustness and further enhancing overall system performance.
6 Generative AI Use Disclosure
Generative AI tools were used for limited language editing purposes, including improving clarity and correcting grammatical issues. No substantive content, analysis, or conclusions were generated by AI. The authors remain fully responsible for the content of this manuscript.